# Pebble Games and Algebraic Proof Systems

### v1

@@ 3.2 Size-degree tradeoffs for MC
[Sav98] There is a family of directed graphs $\{G_{n}\}_{n=0}^{\infty}$ having $\Theta(n^{2})$ vertices each and with $\mathsf{Black}(G_{n})=\Theta(n)$ for which any black pebbling strategy with $\mathsf{Black}(G_{n})$ pebbles requires at least $2^{\Omega(n\log n)}$ steps while there is a pebbling strategy with $\mathsf{Black}(G_{n})+1$ pebbles and $O(n^{2})$ steps.
@@
There is a family of unsatisfiable formulas $\{F_{n}\}_{n=0}^{\infty}$ with $F_{n}$ having $O(n^{2})$ variables and $d_{n}\in O(n)$ such that $F_{n}$ has a MC refutation of degree $d_{n}$ but any MC refutation with this degree requires size $2^{\Omega(n\log n)}.$ On the other side there is a MC refutation of $F_{n}$ with degree $d_{n}+1$ and size O$(n^{3}).$
For a second example we use a robust time-space result from [Nor15].
@@
[Nor12] Let $g(n)$ be any arbitrarily slow growing function $\omega(1)\leq g(n)=O(n^{1/7})$, and let $\epsilon>0$. There is a family of directed graphs $\{G_{n}\}_{n=0}^{\infty}$ having $\Theta(n)$ vertices each and with $\mathsf{Black}(G_{n})=2g(n)+O(1)$, for which there is a black pebbling strategy for $G_{n}$ in time $O(n)$ and space $O(n/g^{2}(n))^{1/3}$ simultaneously. Moreover, any pebbling strategy for $G_{n}$ with space at most $O(n/g^{2}(n))^{1/3-\epsilon}$ requires superpolynomial time in $n$.
@@
4 Pebble Games and Variable Space
The equality between degree and pebbling price for the cases of Monomial Calculus and black pebbling from the previous section, as well as for Nullstellensatz and reversible pebbling from [dRMNR21] cannot be extended to the case of Polynomial Calculus and black-white pebbling price since as already mentioned, it was proven in [BCIP02] that for any DAG $G$, $\deg_{\mathrm{PC}}(\mathrm{Peb}(G))=O(1).$ We show in this section that the correspondence between the three pebbling variations and the proof systems holds if we consider the variables space measure instead.
It can be seen in the proof of Theorem 9, that not only the minimum degree of a monomial calculus refutation of $\mathrm{Peb}(G)$, but also the minimum variable space is bounded by the black pebbling price of $G$. The same can be observed in the proof of Theorem 3.1 in [dRMNR21] for the case of Nullstellensatz and reversible pebbling. Considering the trivial fact that variable space measure is always greater or equal that the degree needed for the refutation of a formula in all three proof systems $\mathrm{NS},\mathrm{MC}$ and $\mathrm{PC}$, and considering Theorems 14 as well as Theorem 3.4 in [dRMNR21] this implies:
Observation 25.
@@ 4 Pebble Games and Variable Space
For the case of black-white pebbling it is known that for the Resolution proof system, the variable space needed in a refutation of $\mathrm{Peb}(G)$ equals $\mathsf{BW}(G).$ The inclusion from left to right is from [Ben02] while the other inclusion appeared in [Her08]. This results can be extended to stronger proof systems using the following result:
@@
[BNT13], [Raz18]. Let $S$ be a proof system that can simulate Resolution step by step without including new variables. For every unsatisfiable formula $F$, $\mathrm{VSpace}_{S}(F\vdash)=\mathrm{VSpace}_{\mathsf{Res}}(F\vdash)$.
@@
Which together with Observation 25 shows the equivalence between variable space in the proof systems and the pebbling price in the three variations of the game.
4.1 Variable Space Separations
These observations allow us to use pebbling results to obtain separation in the variable space complexity in the algebraic proof systems. The reason why these results do not contradict Lemma 26, is that $\mathrm{MC}$ (or $\mathrm{NS})$ cannot simulate Resolution step by step since the intermediate polynomials in the simulation are not necessarily monomials.
For the variable space separations between $\mathrm{NS}$ and $\mathrm{MC}$ on pebbling formulas, the same degree separations given in Subsection 3.1 hold, since as we have seen, for this kind of formulas the variable space and the degree coincide in both proof systems. For the case of $\mathrm{MC}$ versus $\mathrm{PC}$, it is known that for any DAG $G$, the separation between the black and the black-white pebbling prices can be at most quadratic [Mey81]. This limits the variable space gap between $\mathrm{MC}$ and $\mathrm{PC}$ that can be obtained using pebbling formulas. In [Wil88] a family of graphs is given that shows an asymptotic separation between the black-white and black pebbling prices. Translating this to our context we obtain:
@@ 4.1 Variable Space Separations
An optimal quadratic separation between the black-white and black pebbling price was given in [KS91] but for a family of graphs having exponentially many vertices respect to their pebbling price. This implies:
@@
5 Conclusions and Open Questions
We have proven a strong connection between the black pebble game and the Monomial Calculus proof system by showing that the degree and size bounds required simultaneously in a MC refutation of the pebbling formula for a DAG $G$ closely correspond to the number of pebbles and the time in a pebbling strategy for $G$. Together with the connections between the complexity of MC refutations and the Graph Isomorphism formulas form [BG15], these results underline the importance of MC as a natural proof system between NS and PC. Our results improve the known relations between the complexities of pebble games and algebraic proof systems and implies strong degree-size tradeoffs for the $\mathrm{MC}$ system as well as degree separations between $\mathrm{NS},\mathrm{MC}$ and $\mathrm{PC}$.

### v2

@@ 3.2. Size-degree tradeoffs for MC
[31] There is a family of directed graphs $\{G_{n}\}_{n=0}^{\infty}$ having $\Theta(n^{2})$ vertices each and with $\mathsf{Black}(G_{n})=\Theta(n)$ for which any black pebbling strategy with $\mathsf{Black}(G_{n})$ pebbles requires at least $2^{\Omega(n\log n)}$ steps while there is a pebbling strategy with $\mathsf{Black}(G_{n})+1$ pebbles and $O(n^{2})$ steps.
@@
[…]
For a second example we use a robust time-space result from [26].
@@
[24] Let $g(n)$ be any arbitrarily slow growing function $\omega(1)\leq g(n)=O(n^{1/7})$, and let $\epsilon>0$. There is a family of directed graphs $\{G_{n}\}_{n=0}^{\infty}$ having $\Theta(n)$ vertices each and with $\mathsf{Black}(G_{n})=2g(n)+O(1)$, for which there is a black pebbling strategy for $G_{n}$ in time $O(n)$ and space $O(n/g^{2}(n))^{1/3}$ simultaneously. Moreover, any pebbling strategy for $G_{n}$ with space at most $O(n/g^{2}(n))^{1/3-\epsilon}$ requires superpolynomial time in $n$.
@@
4. Pebble Games and Variable Space
The equality between degree and pebbling price for the cases of Monomial Calculus and black pebbling from the previous section, as well as for Nullstellensatz and reversible pebbling from [15] cannot be extended to the case of Polynomial Calculus and black-white pebbling price since as already mentioned, it was proven in [10] that for any DAG $G$, $\deg_{\mathrm{PC}}(\mathrm{Peb}(G))=O(1).$ We show in this section that the correspondence between the three pebbling variations and the proof systems holds if we consider the variables space measure instead.
It can be seen in the proof of Theorem 1, that not only the minimum degree of a monomial calculus refutation of $\mathrm{Peb}(G)$, but also the minimum variable space is bounded by the black pebbling price of $G$. The same can be observed in the proof of Theorem 3.1 in [15] for the case of Nullstellensatz and reversible pebbling. Considering the trivial fact that variable space measure is always greater or equal than the degree needed for the refutation of a formula in all three proof systems $\mathrm{NS},\mathrm{MC}$ and $\mathrm{PC}$, and considering Theorems 4 as well as Theorem 3.4 in [15] this implies:
{obs}
@@ 4. Pebble Games and Variable Space
For the case of black-white pebbling it is known that for the Resolution proof system, the variable space needed in a refutation of $\mathrm{Peb}(G)$ equals $\mathsf{BW}(G).$ The inclusion from left to right is from [7] while the other inclusion appeared in [18]. This results can be extended to stronger proof systems using the following result:
@@
[4], [30]. Let $S$ be a proof system that can simulate Resolution step by step without including new variables. For every unsatisfiable formula $F$, $\mathrm{VSpace}_{S}(F\vdash)=\mathrm{VSpace}_{\mathsf{Res}}(F\vdash)$.
@@
Together with Observation 4 this shows the equivalence between variable space in the proof systems and the pebbling price in the three variations of the game.
4.1. Variable Space Separations
These observations allow us to use pebbling results to obtain separation in the variable space complexity in the algebraic proof systems. The reason why these results do not contradict Lemma 14, is that $\mathrm{MC}$ (and $\mathrm{NS})$ cannot simulate Resolution step by step since the intermediate polynomials in the simulation are not necessarily monomials.
For the variable space separations between $\mathrm{NS}$ and $\mathrm{MC}$ on pebbling formulas, the same degree separations given in Subsection 3.1 hold, since as we have seen, for this kind of formulas the variable space and the degree coincide in both proof systems. For the case of $\mathrm{MC}$ versus $\mathrm{PC}$, it is known that for any DAG $G$, the separation between the black and the black-white pebbling prices can be at most quadratic [22]. This limits the variable space gap between $\mathrm{MC}$ and $\mathrm{PC}$ that can be obtained using pebbling formulas. In [34] a family of graphs is given that shows an asymptotic separation between the black-white and black pebbling prices. Translating this to our context we obtain:
@@ 4.1. Variable Space Separations
An optimal quadratic separation between the black-white and black pebbling price was given in [21] but for a family of graphs having exponentially many vertices respect to their pebbling price. This implies:
@@
5. Conclusions and Open Questions
We have proven a strong connection between the black pebble game and the Monomial Calculus proof system by showing that the degree and size bounds required simultaneously in a MC refutation of the pebbling formula for a DAG $G$ closely correspond to the number of pebbles and the time in a pebbling strategy for $G$. Together with the connections between the complexity of MC refutations and the Graph Isomorphism formulas from [9], these results underline the importance of MC as a natural proof system between NS and PC. Our results improve the known relations between the complexities of pebble games and algebraic proof systems and implies strong degree-size tradeoffs for the $\mathrm{MC}$ system as well as degree separations between $\mathrm{NS},\mathrm{MC}$ and $\mathrm{PC}$.